# Normal Quaternionic Matrices and Finitely Generated Witt Rings

### v1

@@
[…]
Normal Quaternionic Matrices and Finitely Generated Witt Rings
Nico Lorenz N. Lorenz, Fakultät für Mathematik, Ruhr-Universität Bochum, Universitätsstraße 150, 44780 Bochum, Germany nico.lorenz@ruhr-uni-bochum.de  and  Alexander Schönert A. Schönert, Fakultät für Mathematik, Technische Universität Dortmund, Vogelpothsweg 87, 44227 Dortmund Germany alexander.schoenert@tu-dortmund.de
(Date: May 20, 2025)
@@ 0. Introduction
In attempts to characterize (finitely) generated Witt rings over fields of characteristic different from two, there have been abstract approaches for more than 50 years now, starting with the work of Knebusch, Rosenberg, and Ware in the early 1970s ([KnebuschRosenbergWare_WittRings]). In our context, abstract means that we try to grasp the ring structure of Witt rings by mimicking the structure of the square class group along with the represented values of binary quadratic forms, leading to quadratic form schemes as introduced by Cordes in [Cordes_FiniteNoQuaternions], or the behavior of 2-fold Pfister forms and their associated quaternion algebras, leading to quaternionic structures as in Marshall’s book in 1980 [Marshall_AbstractWittRings]. The idea of looking at Witt rings from this perspective has its origin in the following result of Cordes from 1973. Note that we assume all fields to be of characteristic different from two whenever we do not explicitly consider the field $\mathbb{F}_{2}$ with cardinality $2$.
@@
We focus on the approach to describe Witt rings using the properties of quaternion algebras (resp. their norm forms which are $2$-fold Pfister forms) and mainly follow Marshall’s book [Marshall_AbstractWittRings] for our basic notations and definitions.
All so far known examples of finitely generated Witt rings are of the so called elementary type, meaning that they can be derived from certain Witt rings only using two simple construction types. This gives rise to the elementary type conjecture (cf. Remark 2.5) stating that in fact all finitely generated Witt rings are of such type. It is known that the elementary type conjecture holds for Witt rings and thus quaternionic structures over fields $F$ of small order $n:=\operatorname{log}_{2}(|F^{\ast}/F^{\ast 2}|)\in\mathbb{N}_{0}$. The case $n\leq 2$ can be taken care of trivially (cf. Example 1.2), while the case $n=3$ was first done by Kula, Szczepanik and Szymiczek [KulaSzczepanikSzymiczek_8SquareClasses] and the case $n\in\{4,5\}$ by Carson and Marshall [CarsonMarshall_Decomposition]. The case $n=4$ was also dealt with by Szczepanik in 1985 independently ([Szczepanik_RadicalNotExceeding16]). In this paper we introduce a new perspective involving an abstract version of the $2$-torsion part of the Brauer group and developing normal quaternionic matrices which uniquely display quaternionic structures. We apply this to prove some hitherto open cases for the elementary type conjecture, in particular showing that the elementary type conjecture holds for order up to $6$. We also verify the elementary type conjecture for order up to $7$ in all but a few cases in which our computations get too time consuming. While the theoretical parts from Section 1 to Section 5 were developed by the second author in terms of his PhD thesis, the computational approach described in Section 6 was obtained by a joint effort of the authors.
@@ 1. Quaternionic Structures
We summarize the main definitions and results needed for this article from Marshall’s book [Marshall_AbstractWittRings]. The results of classical quadratic form theory mentioned here can, for example, be found in Lam’s standard textbook [Lam_IntroQF].
@@
Let $G$ be a multiplicative group of exponent $2$ with neutral element $1\in G$ and a distinguished element $-1\in G$ (the case $-1=1$ being possible). Write $-a:=(-1)a$ for $a\in G$. Let $Q$ be a set with distinguished element $0\in Q$ and $q:G\times G\to Q$ a surjective map. We call the elements of $Q$ quaternions. We say that $S:=(G,-1,q)$ is a quaternionic structure if the following hold for all $a,b,c,d\in G$:
@@
    Every field $F$ defines a quaternionic structure, using $G_{F}:=F^{\ast}/F^{\ast 2}$ and $q_{F}(a,b):=(a,b)_{F}$, the quaternion algebra with slots $a$ and $b$ over $F$ (for (Q4), see [Lam_IntroQF, Common Slot Theorem III.4.13]). Here, the zero-element in $Q$ can be viewed as the split quaternion algebra or the zero in the Witt ring. We write $S_{F}:=(G_{F},-F^{\ast 2},q_{F})$.
@@
    It is an immediate consequence of the axioms (Q1), (Q2) and (Q3) that any quaternionic structure $(G,-1,q)$ satisfies $q(a,1)=0$ and for all $a,b\in G$, we have:
    $q(a,b)=q(a,-ab)=q(b,-ab)$
@@
  • There are exactly two non isomorphic totally degenerate quaternionic structures of order $n\in\mathbb{N}$, depending on the choice whether or not $-1=1$ in the square class group $G$ and defined by the unique map $G\times G\to\{0\}$. For $n=1$, we denote this structure by $\mathbb{L}_{1,0}$ if $-1=1$ and $\mathbb{L}_{1,1}$ if $1$-1$\not$=1- 1 ≠ 1.
  • There are three quaternionic structures $S=(\{1,a\},-1,q)$ of order $1$: Firstly, if $-1=1$, then $q(1,1)=q(1,a)=q(a,1)=q(a,a)=0$, so $S\cong\mathbb{L}_{1,0}$. Secondly, if $-1=a$ and $S\not\cong\mathbb{L}_{1,1}$, then it can be easily verified that $S\cong S_{\mathbb{R}}=:\mathbb{L}_{1}$. Note that also $\mathbb{L}_{1,0}$ and $\mathbb{L}_{1,1}$ are realized by finite fields of cardinality $\equiv_{4}1$ resp. $\equiv_{4}3$.
  • Let $S=(G,-1,q)$ be a finite nondegenerate quaternionic structure of order $n\geq 3$ with set of quaternions $Q$ with $|Q|=2$, called a local (quaternionic) structure. Then $q$ can be viewed as a nondegenerate symmetric bilinear map $G\times G\to\mathbb{F}_{2}$ (recall that $G$ is an $\mathbb{F}_{2}$ vector space). Conversely, any such nondegenerate symmetric bilinear map defines a local structure. There are exactly two local structures for even $n$, written $\mathbb{L}_{n,0}$ (the case $-1=1$) and $\mathbb{L}_{n,1}$ (the case $1$-1$\not$=1- 1 ≠ 1). Also, there is only one local structure for any odd $n$, written $\mathbb{L}_{n}$. All the local quaternionic structures can be realised by certain field extensions of $\mathbb{Q}_{2}$, the $2$-adic completion of $\mathbb{Q}$, cf. [Marshall_AbstractWittRings, Chapter 5].
@@
We do not give the definition of abstract Witt rings here as we do not need it explicitly. We refer to [Marshall_AbstractWittRings, 4.5] where it is shown that the categories of quaternionic structures and abstract Witt rings as defined there are equivalent. We only outline the notation for quadratic forms and important properties that will be employed here. From a classical standpoint, the quadratic forms considered here correspond to nondegenerate and diagonalized quadratic forms in the classical sense:
@@
  • Isometry of quadratic forms (denoted by $\cong$) is defined inductively as in ([Lam_IntroQF, §I, Exercise 19]) and isotropy derived from it canonically. Isometry defines an equivalence relation (technical proof by Marshall in [Marshall_AbstractWittRings, 2.6]) and isometry is invariant under permutation of tuple entries which means that the following is well defined.
@@
  • The value group of $a\in G$ is defined as the set $V(a):=D_{S}(\langle\!\langle a\rangle\!\rangle)$ and can also be written as $\{b\in G\mid q(a,b)=0\}$. Note that in literature, similar to the ambiguity of the sign convention for Pfister forms, this set is sometimes defined as $V(-a)$.
  • Defining the (abstract) Witt ring $W(S)$ of a quaternionic structure $S=(G,-1,q)$ now works in most parts analogue to the classical case, by establishing Witt cancellation and Witt equivalence in the abstract setting.
@@
($AP_{k}$) If φIk(S){0} is anisotropic, then $\text{If $\varphi\in I^{k}(S)\setminus\{0\}$ is anisotropic, then}\dim(\varphi)\geq 2^{k}$
In the case that $S$ comes from a field, the above holds for arbitrary $k\in\mathbb{N}$, which is known as the Arason-Pfister Hauptsatz [ArasonPfister_Krull]. Generally, $AP_{1}$ and $AP_{2}$ hold in abstract Witt rings (cf. [Marshall_AbstractWittRings, 3.23]) while it is an open question whether or not $AP_{k}$ holds for any $k\geq 3$. Especially $AP_{3}$ plays an important role considering the abstract Brauer group as can be seen later. For a quaternionic structure satisfying $AP_{3}$ we henceforth simply say $AP_{3}$ structure.
@@ 2. The Elementary Type Conjecture
In this section, we recollect all ingredients for the elementary type conjecture as stated by Marshall [Marshall_AbstractWittRings, Chapter 5, Problem 4].
@@
Let $S=(G,-1,q)$, $S_{1}=(G_{1},e_{1},q_{1})$ and $S_{2}=(G_{2},e_{2},q_{2})$ be quaternionic structures with sets of quaternions $Q,Q_{1},Q_{2}$.
@@
If in the above proposition, $S\cong S_{F}$ for a field $F$, then $S_{F}\Delta\cong S_{E}$ for the field $E:=F(\!(x)\!)$ of formal Laurent series over $F$. In 1985, applying quite some valuation theory, Kula [Kula_FieldsQFSchemes] showed that if $S_{1}$ and $S_{2}$ are realized by fields, then so is $S_{1}\times S_{2}$.
@@
  • We have $\mathbb{L}_{1}\times\mathbb{L}_{1}\cong\mathbb{L}_{1}\Delta$. This is in fact the only decomposable (meaning that it can be written as the direct product of two structures of order at least 1) group extension (cf. [Marshall_ETC, p. 8]).
Marshall introduced the following normalized decomposition ([Marshall_AbstractWittRings, 5.12]).
@@
Let $S$ be a finite quaternionic structure with $|\operatorname{rad}(S)|=2^{m}$. Choose $\epsilon=0$ if $-1=1$ in $S$ and $\epsilon=1$ else. There exist indecomposable nondegenerate quaternionic structures $S_{1},\ldots,S_{k}$ such that
@@
Let $S=(G,-1,q)$ be a quaternionic structure. We define the level of $S$ as
@@
where ’torsion form’ means that there exists $n\in\mathbb{N}$ such that $n\times\varphi$ is hyperbolic. In the classical sense, this is often referred to as the general u-invariant (cf. [Lam_IntroQF, Chapter XI]).
@@
$\mathbb{L}_{2k-1}$ $12k-$ 1 4 4 4 $22k-$ 2 $22k-$ 2
$\mathbb{L}_{2k,0}$ $2k$ 1 2 4 $2k$ $12k-$ 1
$\mathbb{L}_{2k,1}$ $2k$ 2 3 4 $12k-$ 1 $12k-$ 1
@@
There have been many attempts to find a field with finitely many square classes and finite level $\geq 8$. This gave rise to the level conjecture (LC) which states that all finite quaternionic structures (coming from fields) have level in $\{1,2,4,\infty\}$. It holds for elementary type structures and thus the LC would follow from the ETC. In fact, the previous example shows that every elementary type structure has level in $\{1,2,4,\infty\}$, Pythagoras number in $\{1,2,3,4,5\}$ and $u$-invariant in $\{0\}\cup\{2^{m}\mid m\in\mathbb{N}_{0}\}$ and all these values can be obtained (though not independently, since for example the Pythagoras number cannot exceed the level by more than $1$). The ’smallest’ example for a quaternionic structure with pythagoras number $5$ is $\mathbb{L}_{3}\Delta$. Note that the LC is known to hold for Witt rings (of fields) of order up to 8 by a result of Becher in 2001 [Becher_NoSquareClasses].
@@
Let $S=(G,-1,q)$ be a quaternionic structure and $k\in\mathbb{N}_{0}$. We say that $a\in G$ is $k$-rigid if $\dim(D_{S}(\langle 1,a\rangle))=k$. If $G$ contains a $k$-rigid element, we say that $S$ is $k$-rigid. If $k\geq 1$ and $G$ additionally does not contain a $(k-1)$-rigid element, we say that $S$ is strictly $k$-rigid. We usually only say rigid and strictly rigid instead of $1$-rigid and strictly $1$-rigid.
Note that $S$ is 00-rigid if and only if $1\in G$ is 00-rigid if and only if $S$ is pythagorean.
@@
    Marshall in [Marshall_AbstractWittRings, 6.23]: If $S$ is pythagorean, then every indecomposable factor of $S$ is isomorphic to $\mathbb{L}_{1}$ or a group extension of $\mathbb{L}_{1}$. In particular, $S$ is of elementary type.
@@
    Kula in [Kula_FinitelyGeneratedWittRings, 3.12]: Every quaternionic structure with at most 12 quaternions is of elementary type.
@@
    Carson and Marshall in [CarsonMarshall_Decomposition, § 5]: If $n\leq 5$, then $S$ is of elementary type.
@@
    Assuming that every quaternionic structure of order up to $n-1$ is of elementary type, $S$ is also of elementary type in the following cases:
@@
    • $S$ is nondegenerate and rigid (Bos in [Bos_StructureWittRing, 2.10])
    • $s(S)=1$ and $S$ is $2$-rigid (Carson and Marshall in [CarsonMarshall_Decomposition, 3.10])
    • $s(S)=2$ and there is a $2$-rigid element in $D_{S}(\langle 1,1\rangle)$ (Cordes in [Cordes_BinaryRepr4, 3.1/4.1])
    • $s(S)=4$ and $1$ is a $2$-rigid element (Cordes in [Cordes_BinaryRepr4, 2.6]).
To close this section, we include a technical result to count elementary types. This is a slightly refined version of the one given in ([CarsonMarshall_Decomposition, 4.5]).
@@
¯$\overline{\mathcal{E}}_{k}(n):=\{S\in\mathcal{E}_{k}(n)\mid S\text{ is indecomposable}\}.$
[…]
@@
      $\overline{e}_{k}(n)=\begin{cases}2,&\text{ if }k+1=n\text{ is even}\\ 1,&\text{ if }k+1=n\text{ is odd}\\ 0,&\text{ else.}\end{cases}$
@@
(a): The degenerate structures are clearly the ones having $k>0$ in the above notation. If $S$ is a degenerate structure of level $1$, it can uniquely be written as $S\cong\mathbb{L}_{1,0}\times S^{\prime}$ for $S^{\prime}\in\mathcal{E}^{\prime}_{n-1}$. This shows that there are $e^{\prime}(n-1)$ degenerate structures of level $1$ and order $n$. If $S$ is a degenerate structure of level different from $1$, we uniquely write $S\cong\mathbb{L}_{1,1}\times S^{\prime\prime}$ for a structure $S^{\prime\prime}$ of order $n-1$ that is either of level different from $1$ or both nondegenerate and of level $1$. Since the number of degenerate structures of level $1$ and order $n-1$ is $e^{\prime}(n-2)$ by the above argument, there are $(e(n-1)-e^{\prime}(n-1))+(e^{\prime}(n-1)-e^{\prime}(n-2))=e(n-1)-e^{\prime}(n-2)$ choices for $S^{\prime\prime}$.
@@
(d) (d)(i): The set $\mathcal{E}_{0}(n)$ contains exactly the pythagorean quaternionic structures of order $n$. Since $n>2$ and local structures are not pythagorean, such a structure is indecomposable if and only if it is the group extension of a pythagorean structure of order $n-1$. Hence, $\overline{e}_{0}(n)=e_{0}(n-1)$. For the second part, we have to count the number of products of indecomposable pythagorean structures of smaller orders that are adding up to $n$. In other words, we just count the number of possible products for each partition $(m_{0,1},\ldots,m_{0,n})\in P_{0,n}$ of the order $n$, we can choose $m_{0,i}$ factors from $\overline{\mathcal{E}_{0}}(i)$ for all $i$. This gives the desired equation.
(d)(ii): The first part follows immediately as the strictly rigid quaternionic structures are exactly the group extensions that are not pythagorean (cf. results stated in Proposition 2.8). For the second part, note that the nondegenerate strictly rigid quaternionic structures are exactly those with one strictly rigid factor in the normalized decomposition while all other factors have to be pythagorean. Hence, such a structure can be uniquely written as the product of an indecomposable strictly rigid structure of order $j\in\{2,\ldots,n\}$ ($j\not=1$ as the factors are nondegenerate) and a pythagorean structure of order $n-j$.
(d)(iii): The indecomposable elementary type structures that have no rigid or 00-rigid elements are exactly the local structures. They have only two quaternions and are nondegenerate which immediately shows that local structures of order $n$ are strictly $(n-1)$-rigid.
@@
Lemma 2.9 enables us to refine the list Carson and Marshall gave in [CarsonMarshall_Decomposition] calculating $e(n)$ and $e^{\prime}(n)$ for small values of $n$, see Table 2. These values will be used later on for the computations in 6. We get the following:

### v2

@@
Abstract.
We present a new approach to verify the Elementary Type Conjecture for abstract Witt rings $(W,G)$ with small $n=\log_{2}(|G|)$. To do so, we make use of an abstract analogue of the 2-torsion part of the Brauer group. This enables us to describe the entire structure of an abstract Witt ring with $|G|=2^{n}$ in terms of a unique $(n\times n)$-matrix. Via computational search, we find all these matrices for $n$ up to $7$, confirming the Elementary Type Conjecture for such $n$.
@@ 0. Introduction
In attempts to characterize (finitely) generated Witt rings over fields of characteristic different from two, there have been abstract approaches for more than 50 years now, starting with the work of Knebusch, Rosenberg, and Ware in the early 1970s ([KRW72]). In our context, abstract means that we try to grasp the ring structure of Witt rings by mimicking the structure of the square class group along with the represented values of binary quadratic forms, leading to quadratic form schemes as introduced by Cordes in [Cor76], or the behavior of 2-fold Pfister forms and their associated quaternion algebras, leading to quaternionic structures as in Marshall’s book in 1980 [Mar80]. The idea of looking at Witt rings from this perspective has its origin in the following result of Cordes from 1973. Note that we assume all fields to be of characteristic different from two whenever we do not explicitly consider the field $\mathbb{F}_{2}$ with exactly two elements.
@@
We focus on the approach to describe Witt rings using the properties of quaternion algebras (resp. their norm forms which are $2$-fold Pfister forms) and mainly follow Marshall’s book [Mar80] for our basic notations and definitions.
All so far known examples of finitely generated Witt rings are of the so called elementary type, meaning that they can be derived from certain Witt rings only using two simple construction types. This gives rise to the elementary type conjecture (cf. Remark 2.5) stating that in fact all finitely generated Witt rings are of such type. It is known that the elementary type conjecture holds for Witt rings and thus quaternionic structures over fields $F$ of small order $n:=\operatorname{log}_{2}(|F^{\ast}/F^{\ast 2}|)\in\mathbb{N}_{0}$. The case $n\leq 2$ can be taken care of trivially (cf. Example 1.2), while the case $n=3$ was first done by Kula, Szczepanik and Szymiczek [KSS79] and the case $n\in\{4,5\}$ by Carson and Marshall [CM82]. The case $n=4$ was also dealt with by Szczepanik in 1985 independently ([Szc85]).
In this paper we introduce a new perspective involving an abstract version of the $2$-torsion part of the Brauer group and developing normal quaternionic matrices which uniquely display quaternionic structures. We apply this to prove that the elementary type conjecture also holds for quaternionic structures of orders $6$ and $7$. While the theoretical parts from Section 1 to Section 5 were developed by the second author in terms of his PhD thesis, the computational approach described in Section 6 was obtained by a joint effort of the authors.
@@ 1. Quaternionic Structures
We summarize the main definitions and results needed for this article from Marshall’s book [Mar80]. The results of classical quadratic form theory mentioned here can, for example, be found in Lam’s standard textbook [Lam05].
@@
Let $G$ be a multiplicative group of exponent $2$ with neutral element $1\in G$ and a distinguished element $1\in G$ (the case 1=1-1=1 being possible). Write a:=(1)a-a:=(-1)a for $a\in G$. Let $Q$ be a set with distinguished element $0\in Q$ and $q:G\times G\to Q$ a surjective map. We call the elements of $Q$ quaternions. We say that S:=(G,1,q)S:=(G,-1,q) is a quaternionic structure if the following hold for all $a,b,c,d\in G$:
@@
    Every field $F$ defines a quaternionic structure, using $G_{F}:=F^{\ast}/F^{\ast 2}$ and $q_{F}(a,b):=(a,b)_{F}$, the quaternion algebra with slots $a$ and $b$ over $F$ (for (Q4), see [Lam05, Common Slot Theorem III.4.13]). Here, the zero-element in $Q$ can be viewed as the split quaternion algebra or the zero in the Witt ring. We write $S_{F}:=(G_{F},-F^{\ast 2},q_{F})$.
@@
    It is an immediate consequence of the axioms (Q1), (Q2) and (Q3) that any quaternionic structure (G,1,q)(G,-1,q) satisfies $q(a,1)=0$ and for all $a,b\in G$, we have:
    q(a,b)=q(a,ab)=q(b,ab)q(a,b)=q(a,-ab)=q(b,-ab)
@@
  • There are exactly two non isomorphic totally degenerate quaternionic structures of order $n\in\mathbb{N}$, depending on the choice whether or not 1=1-1=1 in the square class group $G$ and defined by the unique map $G\times G\to\{0\}$. For $n=1$, we denote this structure by $\mathbb{L}_{1,0}$ if 1=1-1=1 and $\mathbb{L}_{1,1}$ if $1\not=1$.
  • There are three quaternionic structures $S=(\{1,a\},-1,q)$ of order $1$: Firstly, if 1=1-1=1, then $q(1,1)=q(1,a)=q(a,1)=q(a,a)=0$, so $S\cong\mathbb{L}_{1,0}$. Secondly, if 1=a-1=a and $S\not\cong\mathbb{L}_{1,1}$, then it can be easily verified that $S\cong S_{\mathbb{R}}=:\mathbb{L}_{1}$. Note that also $\mathbb{L}_{1,0}$ and $\mathbb{L}_{1,1}$ are realized by finite fields of cardinality $\equiv_{4}1$ resp. $\equiv_{4}3$.
  • Let S=(G,1,q)S=(G,-1,q) be a finite nondegenerate quaternionic structure of order $n\geq 3$ with set of quaternions $Q$ with $|Q|=2$, called a local (quaternionic) structure. Then $q$ can be viewed as a nondegenerate symmetric bilinear map $G\times G\to\mathbb{F}_{2}$ (recall that $G$ is an $\mathbb{F}_{2}$ vector space). Conversely, any such nondegenerate symmetric bilinear map defines a local structure. There are exactly two local structures for even $n$, written $\mathbb{L}_{n,0}$ (the case 1=1-1=1) and $\mathbb{L}_{n,1}$ (the case $1\not=1$). Also, there is only one local structure for any odd $n$, written $\mathbb{L}_{n}$. All the local quaternionic structures can be realised by certain field extensions of $\mathbb{Q}_{2}$, the $2$-adic completion of $\mathbb{Q}$, cf. [Mar80, Chapter 5].
@@
We do not give the definition of abstract Witt rings here as we do not need it explicitly. We refer to [Mar80, 4.5] where it is shown that the categories of quaternionic structures and abstract Witt rings as defined there are equivalent. We only outline the notation for quadratic forms and important properties that will be employed here. From a classical standpoint, the quadratic forms considered here correspond to nondegenerate and diagonalized quadratic forms in the classical sense:
@@
  • Isometry of quadratic forms (denoted by $\cong$) is defined inductively as in ([Lam05, §I, Exercise 19]) and isotropy derived from it canonically. Isometry defines an equivalence relation (technical proof by Marshall in [Mar80, 2.6]) and isometry is invariant under permutation of tuple entries which means that the following is well defined.
@@
  • The value group of $a\in G$ is defined as the set $V(a):=D_{S}(\langle\!\langle a\rangle\!\rangle)$ and can also be written as $\{b\in G\mid q(a,b)=0\}$. Note that in literature, similar to the ambiguity of the sign convention for Pfister forms, this set is sometimes defined as V(a)V(-a).
  • Defining the (abstract) Witt ring $W(S)$ of a quaternionic structure S=(G,1,q)S=(G,-1,q) now works in most parts analogue to the classical case, by establishing Witt cancellation and Witt equivalence in the abstract setting.
@@
($AP_{k}$) $\text{If $\varphi\in I^{k}(S)\setminus\{0\}$ is anisotropic, then }\dim(\varphi)\geq 2^{k}$
In the case that $S$ comes from a field, the above holds for arbitrary $k\in\mathbb{N}$, which is known as the Arason-Pfister Hauptsatz [AP71]. Generally, $AP_{1}$ and $AP_{2}$ hold in abstract Witt rings (cf. [Mar80, 3.23]) while it is an open question whether or not $AP_{k}$ holds for any $k\geq 3$. Especially $AP_{3}$ plays an important role considering the abstract Brauer group as can be seen later. For a quaternionic structure satisfying $AP_{3}$ we henceforth simply say $AP_{3}$ structure.
@@ 2. The Elementary Type Conjecture
In this section, we recollect all ingredients for the elementary type conjecture as stated by Marshall [Mar80, Chapter 5, Problem 4].
@@
Let S=(G,1,q)S=(G,-1,q), $S_{1}=(G_{1},e_{1},q_{1})$ and $S_{2}=(G_{2},e_{2},q_{2})$ be quaternionic structures with sets of quaternions $Q,Q_{1},Q_{2}$.
@@
If in the above proposition, $S\cong S_{F}$ for a field $F$, then $S_{F}\Delta\cong S_{E}$ for the field E:=F((x))E:=F(\!(x)\!) of formal Laurent series over $F$. In 1985, applying quite some valuation theory, Kula [Kul85] showed that if $S_{1}$ and $S_{2}$ are realized by fields, then so is $S_{1}\times S_{2}$.
@@
  • We have $\mathbb{L}_{1}\times\mathbb{L}_{1}\cong\mathbb{L}_{1}\Delta$. This is in fact the only decomposable (meaning that it can be written as the direct product of two structures of order at least 1) group extension (cf. [Mar04, p. 8]).
Marshall introduced the following normalized decomposition ([Mar80, 5.12]).
@@
Let $S$ be a finite quaternionic structure with $|\operatorname{rad}(S)|=2^{m}$. Choose $\epsilon=0$ if 1=1-1=1 in $S$ and $\epsilon=1$ else. There exist indecomposable nondegenerate quaternionic structures $S_{1},\ldots,S_{k}$ such that
@@
Let S=(G,1,q)S=(G,-1,q) be a quaternionic structure. We define the level of $S$ as
@@
where ’torsion form’ means that there exists $n\in\mathbb{N}$ such that $n\times\varphi$ is hyperbolic. In the classical sense, this is often referred to as the general u-invariant (cf. [Lam05, Chapter XI]).
@@
$\mathbb{L}_{2k-1}$ 2k12k-1 4 4 4 2k22k-2 2k22k-2
$\mathbb{L}_{2k,0}$ $2k$ 1 2 4 $2k$ 2k12k-1
$\mathbb{L}_{2k,1}$ $2k$ 2 3 4 2k12k-1 2k12k-1
@@
There have been many attempts to find a field with finitely many square classes and finite level $\geq 8$. This gave rise to the level conjecture (LC) which states that all finite quaternionic structures (coming from fields) have level in $\{1,2,4,\infty\}$. It holds for elementary type structures and thus the LC would follow from the ETC. In fact, the previous example shows that every elementary type structure has level in $\{1,2,4,\infty\}$, Pythagoras number in $\{1,2,3,4,5\}$ and $u$-invariant in $\{0\}\cup\{2^{m}\mid m\in\mathbb{N}_{0}\}$ and all these values can be obtained (though not independently, since for example the Pythagoras number cannot exceed the level by more than $1$). The ’smallest’ example for a quaternionic structure with pythagoras number $5$ is $\mathbb{L}_{3}\Delta$. Note that the LC is known to hold for Witt rings (of fields) of order up to 8 by a result of Becher in 2001 [Bec01].
@@
Let S=(G,1,q)S=(G,-1,q) be a quaternionic structure and $k\in\mathbb{N}_{0}$. We say that $a\in G$ is $k$-rigid if $\dim(D_{S}(\langle 1,a\rangle))=k$. If $G$ contains a $k$-rigid element, we say that $S$ is $k$-rigid. If $k\geq 1$ and $G$ additionally does not contain a (k1)(k-1)-rigid element, we say that $S$ is strictly $k$-rigid. We usually only say rigid and strictly rigid instead of $1$-rigid and strictly $1$-rigid.
Note that $S$ is 0-rigid if and only if $1\in G$ is 0-rigid if and only if $S$ is pythagorean.
@@
    Marshall in [Mar80, 6.23]: If $S$ is pythagorean, then every indecomposable factor of $S$ is isomorphic to $\mathbb{L}_{1}$ or a group extension of $\mathbb{L}_{1}$. In particular, $S$ is of elementary type.
@@
    Kula in [Kul91, 3.12]: Every quaternionic structure with at most 12 quaternions is of elementary type.
@@
    Carson and Marshall in [CM82, § 5]: If $n\leq 5$, then $S$ is of elementary type.
@@
    Assuming that every quaternionic structure of order up to n1n-1 is of elementary type, $S$ is also of elementary type in the following cases:
    1. (i)
@@
    2. (ii)
      $S$ is nondegenerate and rigid (Bos in [Bos89, 2.10])
    3. (iii)
      $s(S)=1$ and $S$ is $2$-rigid (Carson and Marshall in [CM82, 3.10])
    4. (iv)
      $s(S)=2$ and there is a $2$-rigid element in $D_{S}(\langle 1,1\rangle)$ (Cordes in [Cor93, 3.1/4.1])
    5. (v)
      $s(S)=4$ and $1$ is a $2$-rigid element (Cordes in [Cor93, 2.6]).
To close this section, we include a technical result to count elementary types. This is a slightly refined version of the one given in ([CM82, 4.5]).
@@
$\overline{\mathcal{E}}_{k}(n):=\{S\in\mathcal{E}_{k}(n)\mid S\text{ is indecomposable}\}.$
Furthermore, we set $e_{k}(n):=|\mathcal{E}_{k}(n)|$ and $\overline{e}_{k}(n):=|\overline{\mathcal{E}}_{k}(n)|$. Lastly, we write $\mathcal{E}(n)\ /\ \mathcal{E}^{\prime}(n)\ /\ \mathcal{E}_{\text{d}}(n)\ /\ \mathcal{E}_{\Delta}(n)$ for the sets of elementary types of order n/n\ /\ elementary types of order $n$ with level 1/1\ /\ degenerate elementary types of order n/n\ /\ indecomposable group extensions of order $n$ and respectively $e(n),\ e^{\prime}(n),e_{\text{d}}(n),e_{\Delta}(n)$ for their cardinalities.
@@
[…]
@@
(a): The degenerate structures are clearly the ones having $k>0$ in the above notation. If $S$ is a degenerate structure of level $1$, it can uniquely be written as $S\cong\mathbb{L}_{1,0}\times S^{\prime}$ for $S^{\prime}\in\mathcal{E}^{\prime}_{n-1}$. This shows that there are $e^{\prime}(n-1)$ degenerate structures of level $1$ and order $n$. If $S$ is a degenerate structure of level different from $1$, we uniquely write $S\cong\mathbb{L}_{1,1}\times S^{\prime\prime}$ for a structure $S^{\prime\prime}$ of order n1n-1 that is either of level different from $1$ or both nondegenerate and of level $1$. Since the number of degenerate structures of level $1$ and order n1n-1 is $e^{\prime}(n-2)$ by the above argument, there are $(e(n-1)-e^{\prime}(n-1))+(e^{\prime}(n-1)-e^{\prime}(n-2))=e(n-1)-e^{\prime}(n-2)$ choices for $S^{\prime\prime}$.
@@
(d) (d)(i): The set $\mathcal{E}_{0}(n)$ contains exactly the pythagorean quaternionic structures of order $n$. Since $n>2$ and local structures are not pythagorean, such a structure is indecomposable if and only if it is the group extension of a pythagorean structure of order n1n-1. Hence, $\overline{e}_{0}(n)=e_{0}(n-1)$. For the second part, we have to count the number of products of indecomposable pythagorean structures of smaller orders that are adding up to $n$. In other words, we just count the number of possible products for each partition $(m_{0,1},\ldots,m_{0,n})\in P_{0,n}$ of the order $n$, we can choose $m_{0,i}$ factors from $\overline{\mathcal{E}_{0}}(i)$ for all $i$. This gives the desired equation.
(d)(ii): The first part follows immediately as the strictly rigid quaternionic structures are exactly the group extensions that are not pythagorean (cf. results stated in Proposition 2.8). For the second part, note that the nondegenerate strictly rigid quaternionic structures are exactly those with one strictly rigid factor in the normalized decomposition while all other factors have to be pythagorean. Hence, such a structure can be uniquely written as the product of an indecomposable strictly rigid structure of order $j\in\{2,\ldots,n\}$ ($j\not=1$ as the factors are nondegenerate) and a pythagorean structure of order njn-j.
(d)(iii): The indecomposable elementary type structures that have no rigid or 0-rigid elements are exactly the local structures. They have only two quaternions and are nondegenerate which immediately shows that local structures of order $n$ are strictly (n1)(n-1)-rigid.
@@
Lemma 2.9 enables us to refine the list Carson and Marshall gave in [CM82] calculating $e(n)$ and $e^{\prime}(n)$ for small values of $n$, see Table 2. These values will be used later on for the computations in 6. We get the following: